What mass of barium sulfate (233 g/mol) is produced when 125 mL of a 0.150 M solution of barium chloride is mixed with 25 mL of a 0.175 M solution of iron(III) sulfate
Answer #1

3BaCl2+ Fe2(SO4)3?3BaSO4+ 2FeCl3

0.0125 L x 0.150 M = 0.001875 moles barium chloride

0.0025 L x 0.175 M = 0.0004375 moles iron(iii) sulphate

3 moles of BaCl2 reacts with 1 mole of iron(iii) sulphate

=> 0.001875 moles of BaCl2 reacts with (1/3) * 0.001875 = 0.000625 moles of iron(iii) sulphate

so, here the limiting agent is iron(iii) sulphate

so, the 1 mole of iron(iii) sulphate produces 3 moles of barium sulphate

so, 0.0004375 moles iron(iii) sulphate produces 3*0.0004375 = 0.0013125 moles of barium sulphate

as the molecular weight of barium sulphate = 233 g/mole

so the amount of barium sulphate produce = 233 * 0.0013125 = 0.3058 g
Answer #2

Given that Concentration of Bac, 0.150M Volume of BaC,-125mL Concentration of Fe SO0.175M Volume of Fe.SO4-25mL Molar mass of

Hence, for 0.0187 moles of BaCl,, the number of moles of Fe,SO required would be Moles of Fe,SO,-0.0187molx- -0.00623 mol How